OpenStudy (yacoub1993):

Add or subtract as indicated and write the result in standard form. 5i - (-5 - i) A. -5 + 4i B. -5 - 6i C. 5 + 6i D. 5 - 4i

OpenStudy (anonymous):

5i - (-5 - i) is the same as 5i -1(-5 - i) = 5i +5 + i = 5 + 6i